# LtU-ILI: An All-in-One Framework for Implicit Inference in Astrophysics and Cosmology

### v1

@@ 5.4 Multi-round Inference for Gravitational Waves
Figure 8b illustrates the progression of the cumulative posterior probability, as defined in Equation 4, as a function of training time. The noticeable increase in probability mass for both training and validation datasets is attributed to the model’s focus on learning the posterior density near the target observation $\mathcal{P}(\bm{\theta}|\bm{x}=\bm{x}_{0})$. However, an early increase in validation loss during subsequent rounds suggests the potential for overfitting, indicating the model’s limitations in gleaning further insights from the simulation inputs. Sampling the posterior at $\bm{x}=\bm{x_{0}}$ depicted in Figure 8c, reveals that we are able to constrain each of the four parameters around the fiducial point.
@@ 5.5 Inferring Dust Parameters
Dust makes up approximately 1% of the interstellar medium, but reprocesses around 30% of stellar emission (Silva et al.(1998)Silva, Granato, Bressan, and Danese). Understanding the properties of dust and its impact on observables is, therefore, a key goal of galactic and extragalactic astronomy, and a fundamental uncertainty in downstream cosmological inference analyses. Dust leads to an overall reddening of the emission, parametrized by the optical depth and attenuation law; these are analogous to the normalization and slope, particularly where the latter is parametrized as a power law. Below we show how these parameters can be inferred from observational properties using a simple dust model (Trayford et al.(2015)Trayford, Theuns, Bower, Schaye, Furlong, Schaller, Frenk, Crain, Vecchia, and McCarthy) applied to the Simba cosmological hydrodynamic simulation (Davé et al.(2019)Davé, Anglés-Alcázar, Narayanan, Li, Rafieferantsoa, and Appleby) within an SBI framework.
We first briefly describe our forward model for the stellar emission and dust attenuation. We generated spectra for all galaxies with stellar mass $>10^{10}\;\mathrm{M_{\odot}}$ in the Simba simulation using Synthesizer (Lovell et al. in prep.) [^3] https://flaresimulations.github.io/synthesizer. The integrated intrinsic stellar emission of each galaxy was obtained by linking each stellar particle to the BC03 stellar population synthesis models (Bruzual and Charlot(2003)) based on their age and metallicity. We then modelled the dust attenuation as a wavelength-dependent power law, with slope $\alpha$,
@@
Assuming the same underlying stellar emission model, we generated rest frame optical photometry in the $g$ and $r$ bands for all galaxies in Simba for 1000 simulations on a Latin hypercube, with flat priors on the parameters in the range $\alpha\in[0.5,2.0]$, $\tau_{\mathrm{ISM}}\in[0.01,0.5]$ and $\tau_{\mathrm{BC}}\in[0.3,1.5]$. We then take these simulations and measure the $r$-band luminosity function and the $g-r$ color distribution, and use these as our summary statistics. We perform NPE using an MAF architecture; Figure 9 shows the posteriors on a test parameter – data pair. We show results from training on just the luminosity function, just the color, and the two combined. It is clear that the overall normalisation of the luminosity function constrains the ISM optical depth, $\tau_{\mathrm{ISM}}$, whereas the colour constrains the slope of the attenuation law, $\alpha$. When combined, these parameters are simultaneously tightly constrained. The birth cloud attenuation has a more subtle effect on the luminosity function and colour distribution, but the combination leads to tighter constraints.
Figure 9: Dust parameter inference example on the Simba simulation. Left: example posteriors on the ISM and birth cloud optical depth as well as the slope of the attenuation law, from the distribution of $g-r$ colors (blue), the $r$-band magnitude luminosity function (orange), and the combination of both (green). Right: posterior constraints on each parameter for a range of test simulations.
@@ 5.6 The Mass and Energy Loading of Galactic Winds
A major open question in galaxy formation is the role of feedback in regulating star formation and shaping the stellar-to-halo-mass relation. Recently, Carr et al.(2023)Carr, Bryan, Fielding, Pandya, and Somerville and Pandya et al.(2023)Pandya, Fielding, Bryan, Carr, Somerville, Stern, Faucher-Giguère, Hafen, Anglés-Alcázar, and Forbes showed that supernova-driven galactic winds can heat and stir turbulence in the circumgalactic medium of dwarf galaxies and Milky Way-mass halos. This can suppress the cooling and accretion of gas onto galaxies thereby helping to regulate star formation. The strength of this “preventative feedback” is controlled by two wind parameters: the mass loading factor ($\eta_{\rm M}$) and the energy loading factor ($\eta_{\rm E}$). Together, these describe the specific energy of winds and summarize the fraction of mass and energy produced by supernovae that leave the galaxy and become available for large-scale heating.
Here we want to explore the connection between $(\eta_{\rm M},\eta_{\rm E})$ and the stellar-to-halo-mass relation predicted by numerous realizations of the next-generation semi-analytic model sapphire (Pandya et al in prep.). The model used here is more similar to Carr et al.(2023)Carr, Bryan, Fielding, Pandya, and Somerville than Pandya et al.(2023)Pandya, Fielding, Bryan, Carr, Somerville, Stern, Faucher-Giguère, Hafen, Anglés-Alcázar, and Forbes in that it neglects turbulence and assumes purely thermal heating of the circumgalactic medium (CGM) for simplicity. We assume that $\eta_{\rm M}$ and $\eta_{\rm E}$ both follow power law scalings with halo virial velocity $V_{\rm vir}$ (a proxy for halo mass):
@@
Here, $A_{\rm M}$, $\alpha_{\rm M}$, $A_{\rm E}$ and $\alpha_{\rm E}$ are hyperparameters that govern these power law scalings, and we have ignored any redshift dependence. We generate 5,000 model realizations on a Latin hypercube by uniformly sampling combinations of $A_{\rm M}\in[0.01,100]$, $\alpha_{\rm M}\in[-2,2]$, $A_{\rm E}\in[0.01,1]$ and $\alpha_{\rm E}\in[-2,2]$. All other parameters of the model were fixed to reasonable values following Pandya et al.(2023)Pandya, Fielding, Bryan, Carr, Somerville, Stern, Faucher-Giguère, Hafen, Anglés-Alcázar, and Forbes. With these parameters, the system of ordinary differential equations describing the model (Equations 1, 2, and 4-8 of Pandya et al.(2023)Pandya, Fielding, Bryan, Carr, Somerville, Stern, Faucher-Giguère, Hafen, Anglés-Alcázar, and Forbes) is evolved using mass accretion histories of 120 halos randomly selected from five subvolumes of the TNG100 simulation (Pillepich et al.(2018)Pillepich, Springel, Nelson, Genel, Naiman, Pakmor, Hernquist, Torrey, Vogelsberger, Weinberger, and Marinacci). We find that this is a sufficient number of halos to capture variations in the stellar-to-halo-mass relation with model parameters. We restrict to halos with $\log M_{\rm vir}/M_{\odot}=10-12.4$ at $z=0$, spanning the dwarf and Milky Way regime in which supernova feedback is thought to dominate.
Figure 10: Sets of galaxies with properties $\textbf{d}=\{(M_{\rm vir},M_{*})\}_{i=1}^{n_{d}=120}$ are sensitive to sapphire mass and energy loading parameters $\theta=(\eta_{M},\eta_{E})$. Here we utilize the fishnets architecture to aggregate sets of galaxy properties before performing NPE. Contours are shown at the central $68\%$ and $95\%$ confidence intervals for a single data realization (left). Posterior predictive plots show our NPE network is sensitive to both energy loading hyper-parameters but cannot learn the mass loading hyper-parameters with the $M_{*}-M_{\rm vir}$ relation alone.
@@
We use the fishnets architecture by Makinen et al.(2023)Makinen, Alsing, and Wandelt to extract what information the stellar-to-halo-mass relation contains about $\theta=(\eta_{\rm M},\eta_{\rm E})$. The output from each sapphire realization is a set of 120 independently-evolved halos $\textbf{d}=\{(M_{\rm vir},M_{*},\frac{M_{\rm vir}}{M_{*}})\}_{i=1}^{n_{d}=120}$ where $M_{\rm vir}$ and $M_{*}$ are the $z=0$ halo mass and stellar mass, respectively. The likelihood here is a product of each individual halos’ likelihoods, ($\mathcal{L}(\{d_{i}\}|\theta)=\prod^{n_{d}=120}_{i=0}\mathcal{L}(d_{i}|\theta)$, requiring an aggregation over heterogeneous data distributions. Following the optimal fishnets formalism presented in Makinen et al.(2023)Makinen, Alsing, and Wandelt, we embed data into neural score embeddings and Fisher weights, each parameterized by fully-connected networks of size $[128,128,128]$ with LeakyRelu activations before passing the weighted scores to two NPE networks.
@@
We show the resulting posterior for one test simulation in Figure 10. We find that galaxy energy loading factors are extremely sensitive to halo and stellar mass properties (as expected), while mass loading factors prove more difficult to constrain. This is consistent with Carr et al.(2023)Carr, Bryan, Fielding, Pandya, and Somerville who showed that the stellar-to-halo-mass relation is largely insensitive to variations in $\eta_{\rm M}$ but very responsive to changes in $\eta_{\rm E}$. The physical reason for this is that increasing $\eta_{\rm M}$ leads to a higher CGM density and enhanced cooling of gas back into the galaxy without suppressing star formation, thus leading to a similar $M_{*}-M_{\rm vir}$ relation. Instead, increasing $\eta_{\rm E}$ can heat the CGM, prevent gas accretion, reduce star formation and alter the normalization and shape of the $M_{*}-M_{\rm vir}$ relation. In the future, it will be compelling to use fishnets to understand what additional galaxy properties output by sapphire contain the information needed to better learn both $\eta_{\rm M}$ and $\eta_{\rm E}$.
@@ 6 Discussion
Despite the best practices implemented in LtU-ILI, biases may occur when these assumptions are violated. Before applying these approaches to real observational data, it is crucial to understand these failure modes, as well as how to build inference pipelines to be robust to them.
When training data is limited, ILI methods may overfit to a subset of parameter space and suffer from poor out-of-sample performance. This behavior has led to recent discussion as to whether ILI methods are systematically overconfident (Hermans et al.(2022)Hermans, Delaunoy, Rozet, Wehenkel, Begy, and Louppe). PIT tests (§2.4) on independent validation sets are a useful diagnostic tool for detecting overconfidence. For a fixed simulation budget, this problem can be mitigated through the use of ensembling (Hermans et al.(2022)Hermans, Delaunoy, Rozet, Wehenkel, Begy, and Louppe). Ensembling averages the predictions of multiple, independently trained models together, naturally inflating predictive uncertainty to account for epistemic bias (Lakshminarayanan et al.(2017)Lakshminarayanan, Pritzel, and Blundell). Another approach is the frequentist recalibration of error bars (Masserano et al.(2022)Masserano, Dorigo, Izbicki, Kuusela, and Lee).
@@
Model misspecification occurs when the data-generating function for creating training data is poorly representative of physical processes in the real world (Cannon et al.(2022)Cannon, Ward, and Schmon). This can cause an implicit model to interpret the wrong data-parameter connection, resulting in biased predictions. In astrophysics and cosmology, combating model misspecification often requires running more realistic, high-resolution simulators which can be more computationally demanding. In turn, this reduces the number of available simulations, damaging assumption (1). Alternative approaches include increasing the noise model of a given simulation (e.g., Modi et al.(2023)Modi, Pandey, Ho, Hahn, Blancard, and Wandelt), training NDEs with a conservative loss to inflate uncertainties and counteract misspecification (e.g., Kelly et al.(2023)Kelly, Nott, Frazier, Warne, and Drovandi; Huang et al.(2023)Huang, Bharti, Souza, Acerbi, and Kaski), or using latent functions present within intermediate layers of the Bayesian hierarchical model as diagnostics (e.g., Leclercq(2022)). In addition, Bayesian model comparison (Silva et al.(1998)Silva, Granato, Bressan, and Danese; Handley et al.(2015)Handley, Hobson, and Lasenby) is an effective tool for distinguishing between various models conditioned on real data, but also relies on massive simulation budgets. Recent work such as Evidence Networks (Jeffrey and Wandelt(2024)) has shown that this process can be greatly accelerated through the application of neural networks.
@@
Lastly, implicit inference can be very challenging in the high-dimensional parameter regime (i.e., when $\operatorname{dim}(\bm{\theta})\gg 100$). As the dimensionality of $\bm{\theta}$ increases, the number of parameters necessary to specify a highly-correlated posterior increases exponentially. This problem is of particular relevance in astrophysics and cosmology, in which we often would like to infer voxelized 2D or 3D fields (e.g., Jasche and Wandelt(2013); de Andres et al.(2024)de Andres, Cui, Yepes, De Petris, Ferragamo, De Luca, Aversano, and Rennehan). Recent empirical evidence suggests that Moment Networks (Jeffrey and Wandelt(2020)) and diffusion-based generative models (Song et al.(2020)Song, Sohl-Dickstein, Kingma, Kumar, Ermon, and Poole; Legin et al.(2024)Legin, Ho, Lemos, Perreault-Levasseur, Ho, Hezaveh, and Wandelt) are incredibly effective neural architectures for these tasks. Though LtU-ILI does not currently include these models, we intend to implement them in future work.

### v2

@@ 5.4 Multi-round Inference for Gravitational Waves
Figure 8b illustrates the progression of the cumulative posterior probability, as defined in Equation 4, as a function of training time. The noticeable increase in probability density for both training and validation datasets is attributed to the model’s focus on learning the posterior density near the target observation $\mathcal{P}(\bm{\theta}|\bm{x}=\bm{x}_{0})$. However, an early increase in validation loss during subsequent rounds suggests the potential for overfitting, indicating the model’s limitations in gleaning further insights from the simulation inputs. Sampling the posterior at $\bm{x}=\bm{x_{0}}$ depicted in Figure 8c, reveals that we are able to constrain each of the four parameters around the fiducial point. For this particular GW example, using LtU-ILI allows us to parameterize the simulator (waveform approximant, number of detectors considered, geometry of the detection, duration of the signal [^3] Note, the frequency resolution is only determined by the time duration of the signal, whereas the sampling rate only gives the Nyquist frequency, …), the architectures to train and the metrics to compute with one single set of parameter files, thus making the organization and replication of unified inference tests easier for this particularly difficult problem.
@@ 5.5 Inferring Dust Parameters
LtU-ILI is a practical tool for exploratory discovery, especially useful for investigating novel datasets within which likelihoods are not well-defined. In this application, we study how different observables can be used to constrain different degeneracies within galactic dust models, and how their observational combinations can be used to maximally extract information within an ILI model. Dust makes up approximately 1% of the interstellar medium, but reprocesses around 30% of stellar emission (Silva et al., 1998). Understanding the properties of dust and its impact on observables is, therefore, a key goal of galactic and extragalactic astronomy, and a fundamental uncertainty in downstream cosmological inference analyses (Crocce et al., 2016). Dust leads to an overall reddening of the emission, parameterized by the optical depth and attenuation law; these are analogous to the normalization and slope, particularly where the latter is parameterized as a power law. Below we show how these parameters can be inferred from observational properties using a simple dust model (Trayford et al., 2015) applied to the Simba cosmological hydrodynamic simulation (Davé et al., 2019) within an ILI framework.
We first briefly describe our forward model for the stellar emission and dust attenuation. We generated spectra for all galaxies with stellar mass $>10^{10}\;\mathrm{M_{\odot}}$ in the Simba simulation using Synthesizer (Lovell et al. in prep.) [^4] https://flaresimulations.github.io/synthesizer. The integrated intrinsic stellar emission of each galaxy was obtained by linking each stellar particle to the BC03 stellar population synthesis models (Bruzual and Charlot, 2003) based on their age and metallicity. We then modelled the dust attenuation as a wavelength-dependent power law, with slope $\alpha$,
@@
Assuming the same underlying stellar emission model, we generated rest frame optical photometry in the $g$ and $r$ bands for all galaxies in Simba for 1000 simulations on a Latin hypercube, with flat priors on the parameters in the range $\alpha\in[0.5,2.0]$, $\tau_{\mathrm{ISM}}\in[0.01,0.5]$ and $\tau_{\mathrm{BC}}\in[0.3,1.5]$. We then take these simulations and measure the $r$-band luminosity function and the $g-r$ color distribution, and use these as our summary statistics. We perform NPE using an ensemble of a MAF with 50 hidden features and five neural transformations and an MDN with 50 hidden features and five mixture components.
Figure 9 shows the posteriors on a test parameter – data pair. We show results from training on just the luminosity function, just the color, and the two combined. It is clear that the overall normalisation of the luminosity function constrains the ISM optical depth, $\tau_{\mathrm{ISM}}$, whereas the colour constrains the slope of the attenuation law, $\alpha$. When combined, these parameters are simultaneously tightly constrained. The birth cloud attenuation has a more subtle effect on the luminosity function and colour distribution, but the combination leads to tighter constraints. LtU-ILI greatly simplifies the setup of the inference pipelines used here, for rapid testing and comparison of combinations of data sources.
Figure 9: Dust parameter inference example on the Simba simulation. Example posteriors on the ISM and birth cloud optical depth as well as the slope of the attenuation law, from the distribution of $g-r$ colors (blue), the $r$-band magnitude luminosity function (orange), and the combination of both (green).
@@ 5.6 The Mass and Energy Loading of Galactic Winds
Figure 10: Inference of mass and energy loading parameters of sapphire semi-analytic model from sets of galaxy virial and stellar masses, i.e., $\textbf{x}=\{(M_{\rm vir},M_{*},\frac{M_{\rm vir}}{M_{*}})\}_{i=1}^{n_{x}}$. The left plots show contours of an example inferred posterior, while the right plots show the true values vs. predictions across the full test set. Here, we utilize the fishnets architecture to aggregate sets of galaxy properties before performing NPE. Posterior predictive plots show our NPE network is sensitive to both energy loading hyper-parameters but cannot learn the mass loading hyper-parameters with the $M_{*}-M_{\rm vir}$ relation alone.
Lastly, we demonstrate how using custom embedding networks within LtU-ILI can aid posterior recovery for low signal-to-noise problems such as inferring parameters of semi-analytic models of galaxy formation. One such example is studying the poorly constrained role of feedback in regulating star formation and shaping the stellar-to-halo-mass relation. Recently, Carr et al. (2023) and Pandya et al. (2023) showed that supernova-driven galactic winds can heat and stir turbulence in the circumgalactic medium of dwarf galaxies and Milky Way-mass halos. This can suppress the cooling and accretion of gas onto galaxies thereby helping to regulate star formation. The strength of this “preventative feedback” is controlled by two wind parameters: the mass loading factor ($\eta_{\rm M}$) and the energy loading factor ($\eta_{\rm E}$). Together, these describe the specific energy of winds and summarize the fraction of mass and energy produced by supernovae that leave the galaxy and become available for large-scale heating.
Here we want to explore the connection between $\eta_{\rm M}$, $\eta_{\rm E})$, and the stellar-to-halo-mass relation predicted by numerous realizations of the next-generation semi-analytic model sapphire (Pandya et al in prep.). The model used here is more similar to Carr et al. (2023) than Pandya et al. (2023) in that it neglects turbulence and assumes purely thermal heating of the circumgalactic medium (CGM) for simplicity. We assume that $\eta_{\rm M}$ and $\eta_{\rm E}$ both follow power law scalings with halo virial velocity $V_{\rm vir}$ (a proxy for halo mass):
@@
Here, $A_{\rm M}$, $\alpha_{\rm M}$, $A_{\rm E}$ and $\alpha_{\rm E}$ are hyperparameters that govern these power law scalings, and we have ignored any redshift dependence. We generate 5,000 model realizations on a Latin hypercube by uniformly sampling combinations of $A_{\rm M}\in[0.01,100]$, $\alpha_{\rm M}\in[-2,2]$, $A_{\rm E}\in[0.01,1]$ and $\alpha_{\rm E}\in[-2,2]$. All other parameters of the model were fixed to reasonable values following Pandya et al. (2023). With these parameters, the system of ordinary differential equations describing the model (Equations 1, 2, and 4-8 of Pandya et al. 2023) is evolved using mass accretion histories of 120 halos randomly selected from five subvolumes of the TNG100 simulation (Pillepich et al., 2018). We find that this is a sufficient number of halos to capture variations in the stellar-to-halo-mass relation with model parameters. We restrict to halos with $\log M_{\rm vir}/M_{\odot}=10-12.4$ at $z=0$, spanning the dwarf and Milky Way regime in which supernova feedback is thought to dominate.
@@
As an embedding network, we use the fishnets architecture by Makinen et al. (2023) to extract what information the stellar-to-halo-mass relation contains about $\theta=(A_{\rm M},\alpha_{\rm M},A_{\rm E},\alpha_{\rm E})$. The output from each sapphire realization is a set of $n_{x}=120$ independently-evolved galaxies, i.e., $\textbf{x}=\{(M_{\rm vir},M_{*},\frac{M_{\rm vir}}{M_{*}})\}_{i=1}^{n_{x}}$ where $M_{\rm vir}$ and $M_{*}$ are the $z=0$ halo mass and stellar mass, respectively. The likelihood here is a product of each individual galaxies’ likelihoods, ($\mathcal{L}(\{x_{i}\}|\theta)=\prod^{n_{x}}_{i=0}\mathcal{L}(x_{i}|\theta)$, requiring an aggregation over heterogeneous data distributions. Following the optimal fishnets formalism presented in Makinen et al. (2023), we embed data into neural score embeddings and Fisher weights, each parameterized by fully-connected networks of size $[128,128,128]$ with LeakyRelu activations before passing the weighted scores to two NPE networks, a Neural Spline Flow (Durkan et al., 2019) and a Gaussianization Flow (Meng et al., 2020).
@@
We show the resulting true parameter versus predicted posterior plots in Figure 10. We find that galaxy energy loading factors are sensitive to halo and stellar mass properties as expected, while mass loading factors prove more difficult to constrain. This is consistent with Carr et al. (2023) who showed that the stellar-to-halo-mass relation is largely insensitive to variations in $\eta_{\rm M}$ but very responsive to changes in $\eta_{\rm E}$. The physical reason for this is that increasing $\eta_{\rm M}$ leads to a higher CGM density and enhanced cooling of gas back into the galaxy without suppressing star formation, thus leading to a similar $M_{*}-M_{\rm vir}$ relation. Instead, increasing $\eta_{\rm E}$ can heat the CGM, prevent gas accretion, reduce star formation and alter the normalization and shape of the $M_{*}-M_{\rm vir}$ relation. Interestingly, for sapphire realizations run at $A_{\rm M}\lesssim 0.25$ there is some information available to constrain mass-loading parameters, which isn’t accesible for high $A_{\rm M}$. LtU-ILI is able to capture this effect and carefully estimate posterior error bars, demonstrating it’s ability to calibrate inference both in low and high signal-to-noise regimes. In the future, it will be compelling to use fishnets to understand what additional galaxy properties output by sapphire contain the information needed to better learn both $\eta_{\rm M}$ and $\eta_{\rm E}$.
@@ 6 Discussion
Despite the best practices implemented in LtU-ILI, biases may occur when these assumptions are violated. Before applying these approaches to real observational data, it is crucial to understand these failure modes, as well as how to build inference pipelines to be robust to them. Below, we review important considerations provide qualitative guidance for using ILI in observational settings.
@@
Firstly, model misspecification occurs when the data-generating function for creating training data is poorly representative of physical processes in the real world (Cannon et al., 2022). This can cause an implicit model to interpret the wrong data-parameter connection, resulting in biased predictions. In astrophysics and cosmology, combating model misspecification often requires running more realistic, high-resolution simulators which can be more computationally demanding. In turn, this reduces the number of available simulations, damaging assumption (1). Alternative approaches include increasing the noise model of a given simulation (e.g., Modi et al., 2023), training NDEs with a conservative loss to inflate uncertainties and counteract misspecification (e.g., Kelly et al., 2023; Huang et al., 2023), or using latent functions present within intermediate layers of the Bayesian hierarchical model as diagnostics (e.g., Leclercq, 2022).
A more direct approach to addressing model misspecification is through Bayesian model comparison (Silva et al., 1998; Handley et al., 2015). Given observational data $\bm{x}_{o}$, this technique seeks to quantify whether a proposed data-generating model $\mathcal{M}(\bm{\theta})$ is reflective of reality. This is often done through calculation of the model evidence, $p\left(\bm{x}_{o}|\mathcal{M}\right)$, which, given an explicit likelihood model $\mathcal{L}(\bm{x}|\bm{\theta},\mathcal{M})$, can be done numerically, though requiring massive simulation budgets. Recent ILI approaches using the harmonic mean estimator (McEwen et al., 2021; Spurio Mancini et al., 2023) or Evidence Networks (Jeffrey and Wandelt, 2024) have been shown to greatly accelerate these calculations. They also allow us to learn distributions implicitly from simulations, eliminating the need for explicit likelihoods. These novel evidence estimators are a natural addition to LtU-ILI and are planned for a future release. When implemented, these techniques will allow users to form quantitative arguments to addressing model misspecification.
Secondly, even for well-specified models, ILI methods may overfit and suffer from poor out-of-sample performance when training data is limited. This behavior has led to recent discussion as to whether ILI methods are systematically overconfident (Hermans et al., 2022). PIT tests (Section 2.5) on independent validation sets are a useful diagnostic tool for detecting overconfidence. For a fixed simulation budget, this problem can be mitigated through the use of ensembling (Hermans et al., 2022). Ensembling averages the predictions of multiple, independently trained models together, naturally inflating predictive uncertainty to account for epistemic bias (Lakshminarayanan et al., 2017). Alternatively, approximate Bayesian neural networks (Blundell et al., 2015; Gal and Ghahramani, 2016) seek to constrain posteriors over model parameters during training, which they then subsequently propagate to model outputs to increase posterior variance. Lastly, once a model has been trained, its predictions can be reweighted through frequentist recalibration (Masserano et al., 2022) which guarantees accurate posterior coverage over held-out test data. While approximate Bayesian networks and frequentist recalibration are not yet implemented in LtU-ILI, their integration is considered for future versions of the codebase.
Overfitting can also be calibrated through careful hyperparameter selection. Choices of input data, neural architectures, and training procedures all play a quantifiable role in the final accuracy and out-of-sample bias measured in an ILI model. Oftentimes, selection of hyperparameters is a trade-off between choosing a flexible, expressive model which produces high-accuracy but is prone to overfitting, against a shallow, weaker model with more generalizable predictions. This is commonly referred to as the ‘bias-variance tradeoff’ (Belkin et al., 2019), and can be interpreted as placing assumptions (1) and (3) in direct tension. To address this, practitioners often perform automated searches over hyperparameters to find configurations which minimize variance within a given bias tolerance. Often, these searches are accelerated through the use of Bayesian optimization tools such as optuna (Akiba et al., 2019). These can be used to motivate ILI design choices directly from data, a necessary step for scientific studies. In the LtU-ILI repository, we provide examples of hyperparameter optimization through grid searching, and we plan to link LtU-ILI to optimal search tools such as optuna in future iterations.
@@
Lastly, implicit inference can be very challenging in the high-dimensional parameter regime (i.e., when $\operatorname{dim}(\bm{\theta})\gg 100$). As the dimensionality of $\bm{\theta}$ increases, the number of parameters necessary to specify a highly-correlated posterior increases exponentially. This problem is of particular relevance in astrophysics and cosmology, in which we often would like to infer voxelized 2D or 3D fields (e.g., Jasche and Wandelt, 2013; de Andres et al., 2024). Recent empirical evidence suggests that Moment Networks (Jeffrey and Wandelt, 2020) and diffusion-based generative models (Song et al., 2020; Legin et al., 2024) are remarkably effective neural architectures for these tasks. Though LtU-ILI does not currently include these models, we intend to implement them in future work.